Test the hypotheses, given p=0.42, p is not equal to 0.42, sample size=150, sample proportion=0.45, alpha =0.05


1
Expert's answer
2022-05-04T13:23:42-0400

The following null and alternative hypotheses for the population proportion needs to be tested:

"p=0.42"

"p\not=0.42"

This corresponds to a two-tailed test, for which a z-test for one population proportion will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a two-tailed test is "z_c = 1.96."

The rejection region for this two-tailed test is "R = \{z: |z| > 1.96\}

."

The z-statistic is computed as follows:


"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}=\dfrac{0.45-0.42}{\sqrt{\dfrac{0.42(1-0.42)}{150}}}"

"\approx0.7444"

Since it is observed that "|z| = 0.744 \le1.96= z_c ," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value is "p =2 P(Z>1.7444)=0.456635," and since "p = 0.4566 35>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population proportion "p" is different than 0.42, at the "\alpha = 0.05" significance level.